OpenStudy (anonymous):

Rewrite the equation y = 3x^2 + 6x + 2 in the alternate form and state the vertex point.

jimthompson5910 (jim_thompson5910):

do you see how y = 3x^2 + 6x + 2 is in the form y = ax^2 + bx + c ?

OpenStudy (anonymous):

a = 3 b = 6 c = 2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

now plug the values of a,b into h = -b/(2a) what is the value of h?

OpenStudy (anonymous):

h = -1?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now plug that value in for x to find y y = 3x^2 + 6x + 2 y = 3(-1)^2 + 6(-1) + 2 y = ???

OpenStudy (anonymous):

Is it -1?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so k = -1

jimthompson5910 (jim_thompson5910):

we have a = 3, h = -1 and k = -1 plug those into y = a(x-h)^2 + k

OpenStudy (anonymous):

Does x = 6?

jimthompson5910 (jim_thompson5910):

where are you getting x = 6 ?

OpenStudy (anonymous):

Oh thought x = 6 from the problem. Dont know what x is.

jimthompson5910 (jim_thompson5910):

x is just a variable to be left alone really

jimthompson5910 (jim_thompson5910):

what did you get when you plugged those values into y = a(x-h)^2 + k

OpenStudy (anonymous):

Is it 11?

jimthompson5910 (jim_thompson5910):

don't replace x though

OpenStudy (anonymous):

Oh keep x as just x?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

So is this right? y - 3(x - (-1)^2 + - 1 . I think I set it up wrong.

OpenStudy (anonymous):

I meant y =

jimthompson5910 (jim_thompson5910):

if you simplified that, you'd get what?

OpenStudy (anonymous):

Does the 3 go with the things inside the ( )

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):